OpenStudy (anonymous):

Write the explicit formula for the geometric sequence. Then find the fifth term in the sequence. \(a_1 = –4,~a_2 = 8,~a_3 = –16\) A) \(a_n = –4 · (2)^n; –64\) B) \(a_n = –4 · (–2)^{n–1}; –64\) C) \(a_n = –4 · (–2)^n; 128\) D) \(a_n = –2 · (–4)^{n–1}; –512\)

OpenStudy (campbell_st):

well the general term in a geometric sequence is \[a_{n} =a_{1} \times r^{n -1}\] you know the 1st term.... you need to common ratio r..... before you can find the 5th term hope it helps

OpenStudy (anonymous):

Using what you said would B be correct? \(a_{n} =a_{1} \times r^{n -1}\) \(a_{n} =-4 \times -2^{n -1}\)

OpenStudy (campbell_st):

yeap...thats it, you now just need to substitute n = 5 and then evaluate